# First-Principles Study of the Fermi Surface Topology of CeCu2Si2

### v1

@@ 2.3 Fermi Surface Topology
Analysis of the Fermi surface topology gives a clear indication of the impact of the CFS in CeCu2Si2. There are numerous theoretical studies and experimental measurements that look at the Fermi surface. Based on the combined efforts of magnetic oscillation studies [5] and ARPES experiments[28], it has been experimentally confirmed that there are two sheets in the Fermi surface. Theoretical methodologies from the LDA[6, 7] to LDA+ in the random phase approximation (RPA)[24, 18, 19, 28, 27] to renormalized band theory[8, 17] and DMFT[16] have been applied to understand the Fermi surface topology. The theoretical results appear to depend heavily on the method, choice of electronic correlation strength, and the presence of CFS.
@@
In the early works of Harima et al.[6, 7], the LDA method predicted three Fermi surface sheets. Our GGA calculations at , agree closely with both the qualitative character of the Fermi sheet topologies as well as the dHvA simulation results for the maximal frequencies and cyclotron masses. In our survey of the Hubbard parameter strength in the GWA[35], we have tracked the evolution of the Fermi surface topology (FIG. 3), and simulated the dHvA extremal frequencies for an external B field parallel to the -axis. The dHvA results for the Fermi surface obtained using the simulation method of Rourke and Julian[41] described in section 4.2 are given in Table 2, which shows the results when CFS is included and when it is neglected. For , the impact of CFS on the Fermi surface topology is negligible.
@@
It is well documented that better agreement with the current experimental consensus can be achieved in the Fermi Surface by including strong electron-electron correlations (via a Coulomb interaction). The DFT++RPA type methods tend to over estimate the impact of the 4f electrons in the conduction bands, but they are capable of giving results that compare qualitatively well with experiment. These methods also appear to be successful with smaller values according to Ikeda et al.[18]. In our calculations using a GWA approach without CFS the correlation strength needs to be large in order to predict a Fermi surface containing two sheets, rather than three. Even at eV this method does not give the expected cylindrical heavy band predicted by the DFT++RPA and renormalized band theory methods (FIG. 3). Additionally, we explored a DMFT approach (neglecting CFS) discussed in the Supplementary Materials[35]. These Fermi surface topology results compare nearly exactly to our GWA calculation at the same value of , which is shown in FIG.S10. There have been several DMFT studies of CeCu2Si2[16, 20, 25]. The study by Pourovskii et al..[16] included CFS and predicted a single sheet in the Fermi surface, which is at odds with previous DFT+ methods and experimental measurement. It is, however in close agreement with our GWA+CFS results at eV (FIG. 3).
@@
The electronic structure results from section 2.2 are an indication that the CFS effects are significant to the band dispersion and therefor the Fermi surface. This is demonstrated by the comparison of FIG.3 (a) and (b), where (b) shows the evolution of the Fermi surface topology with increasing with CFS included. At eV with CFS, there are two sheets making up the Fermi surface and the dHvA data is given in Table 2. The qualitative Fermi surface is in close agreement with the renormalized band theory studies of Zwicknagl et al.[8, 17], as well as the DFT++RPA methods[18, 19]. The effective cyclotron mass (eq. 6) calculated with CFS also agrees closely with the renormalized band theory calculations which predicted an effective mass of for the light quasiparticle sheet and for the heavy sheet. Our calculations predict 4.35 and 488 for the light and heavy sheets respectively (see section III.B for the Suplementary Material[35]).
@@ 3 Conclusion
In the last years there have been a number of studies using various methods to describe the Fermi surface of CeCu2Si2. Experimental measurements have confirmed the existence of a Fermi Surface comprised of two sheets[5, 28], several theoretical and experimental studies have explored the importance of the CFS and electron correlation effects on the overall electronic structure and the superconducting pairing[9, 12, 41, 30, 18, 19, 20, 22], and new advances have made it possible to perform more accurate and detailed analysis of the intricate coupling and correlation effects in this complex material.
@@
In this study we have presented a systematic analysis of the electronic structure and Fermi surface topology of CeCu2Si2 as a function of the Coulomb interaction strength with a focus on the CFS effect. We have performed calculations using the GWA to show that the resulting Fermi surface topology is highly sensitive to methodology, electron-electron interactions, and CFS. De Haas van Alphen simulations were applied to the Fermi surfaces to analyze the Fermi sheets. Our GWA results calculated using a Hubbard parameter of eV and including CFS are found to be in agreement with both experimental measurements as well as the renomalized band theory. More broadly, our results demonstrate that the CFS and electronic correlations are on the same order of importance in the electronic structure of CeCu2Si2. The qualitative and quantitative aspect of the Fermi surface are also extremely sensitive to small changes in the electron correlation, particularly in the presence of CFS. Based on the analysis of the electronic DOS, this is due to the complexity of the 4f electron behavior present in the heavy Fermi sheet.
@@
Our results highlight the necessity for a thorough treatment of the electronic coupling in order to better understand the nature of the heavy-fermion superconductivity and formation of Cooper pairs. Future studies of the temperature dependence (which has been touched on in Luo et al..[25]) and possible electron-phonon coupling contribution would aid in forming an even more comprehensive picture of CeCu2Si2.
4 Methods
4.1 Computational Methods
@@
To account for the strong electronic coupling in CeCu2Si2 we combine standard density functional theory with the GWA method implemented in the CyGutz code [32, 38]. This method uses the full-potential, linearly augmented plane wave (FPLAPW) DFT of WIEN2k [42] as its basis and implements a combination of the GWA and the Slave Boson method to account for strong electronic correlations. Relativistic spin-orbit coupling (SOC) effects were included with a k-point grid of and muffin tin radii of 2.50 for the Ce and Cu atoms and 1.93 for the Si atoms, where is the Bohr radius. The cut-off parameter was set to . A generalized gradient approximation (GGA-PBE)[43] exchange correlation functional was used throughout all calculations. The Coulomb interaction strength was tested at values from 0 eV to 6.0 eV[35]. We tested the GWA band renormalization implementation with and without CFS effects included beyond the initial DFT. For all calculations where , Hund’s exchange coupling parameter, , was set at 0.7 eV. For comparison, a GGA+DMFT calculation was performed for eV and eV. Results and details of the methodology are given in the Supplementary Materials, section II[35].
4.2 de Haas van Alphen simulations
@@
Analysis of the Fermi surfaces was done using numerical calculations of the dHvA effect as implemented by Rourke and Julian[41]. By applying a magnetic field to the system, oscillations in the magnetic susceptibility can be determined from the changes in the number of occupied Landau levels as a function of the reciprocal magnetic field, [41, 44]. Then the dHvA frequency can be expressed as
@@ Acknowledgements
We thank Chao Cao, Bogyu Jang, E. M. Nica, F. Steglich and G. Zwicknagl for useful discussions. This work was carried out under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration under Contract No. 89233218CNA000001. The Fermi-surface topology analysis work was partly supported by the LANL LDRD Program and the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. The GWA simulation work was supported by the NNSA Advanced Simulation and Computing Program. It was, in part, supported by the Center for Integrated Nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resource as well as NERSC.The work at Rice has primarily been supported by the National Science Foundation under Grant No. DMR-2220603, by the Robert A. Welch Foundation Grant No. C-1411 and the Vannevar Bush Faculty Fellowship ONR-VB N00014-23-1-2870. Q. S. acknowledges the hospitality and support by a Ulam Scholarship from the Center for Nonlinear Studies at Los Alamos National Laboratory during the initial stage of the work.

### v2

@@ II.3 Fermi Surface Topology
Analysis of the Fermi surface topology gives a clear indication of the impact of the CFS in the presence of strong electron correlations. There are numerous theoretical studies and experimental measurements that look at the Fermi surface of CeCu2Si2. Based on the combined efforts of magnetic oscillation studies [1990Hunt] and ARPES experiments[2021Wu], it has been experimentally confirmed that there are two sheets in the Fermi surface. Theoretical methodologies from the LDA[1991Harima, 1992Harima] to LDA+ in the random phase approximation (RPA)[2008Eremin, 2015Ikeda, 2021Song, 2021Wu, 2024Zhao] to renormalized band theory[1993Zwicknagl, 2016Zwicknagl] and DMFT[2014Pourovskii] have been applied to understand the Fermi surface topology. The theoretical results appear to depend heavily on the method, choice of electronic correlation strength, and the presence of CFS.
@@
In the early works of Harima et al.[1991Harima, 1992Harima], the LDA method predicted three Fermi surface sheets. Our GGA calculations at , agree closely with both the qualitative character of the Fermi sheet topologies as well as the dHvA simulation results for the maximal frequencies and cyclotron masses. In our survey of the Hubbard parameter strength in the GWA[SupMat], we have tracked the evolution of the Fermi surface topology (FIG. 3), and simulated the dHvA extremal frequencies for an external B field parallel to the -axis. The dHvA results for the Fermi surface obtained using the simulation method of Rourke and Julian[2012Rourke] described in section IV.2 are given in Table 2, which shows the results when CFS is included and when it is neglected. For , the impact of CFS on the Fermi surface topology is negligible.
@@
It is well documented that better agreement with the current experimental consensus can be achieved in the Fermi Surface by including strong electron-electron correlations (via a Coulomb interaction). The DFT++RPA type methods tend to over estimate the impact of the 4f electrons in the conduction bands, but they are capable of giving results that compare qualitatively well with experiment. These methods also appear to be successful with smaller values according to Ikeda et al.[2015Ikeda]. In our calculations using a GWA approach without CFS the correlation strength needs to be large in order to predict a Fermi surface containing two sheets, rather than three. Even at eV this method does not give the expected cylindrical heavy band predicted by the DFT++RPA and renormalized band theory methods (FIG. 3). Additionally, we explored a DMFT approach (neglecting CFS) discussed in the Supplementary Materials[SupMat]. These Fermi surface topology results compare nearly exactly to our GWA calculation at the same value of , which is shown in FIG.S10. There have been several DMFT studies of CeCu2Si2[2014Pourovskii, 2020Li, 2020Luo]. The study by Pourovskii et al..[2014Pourovskii] included CFS and predicted a single sheet in the Fermi surface, which is at odds with previous DFT+ methods and experimental measurement. It is, however in close agreement with our GWA+CFS results at eV (FIG. 3).
@@
The electronic structure results from section II.2 are an indication that the CFS effects are significant to the band dispersion and therefor the Fermi surface. This is demonstrated by the comparison of FIG.3 (a) and (b), where (b) shows the evolution of the Fermi surface topology with increasing with CFS included. At eV with CFS, there are two sheets making up the Fermi surface and the dHvA data is given in Table 2. The qualitative Fermi surface is in close agreement with the renormalized band theory studies of Zwicknagl et al.[1993Zwicknagl, 2016Zwicknagl], as well as the DFT++RPA methods[2015Ikeda, 2021Song]. The effective cyclotron mass (eq. 5) calculated with CFS also agrees closely with the renormalized band theory calculations which predicted an effective mass of for the light quasiparticle sheet and for the heavy sheet. Our calculations (with CFS included) predict 4.35 and 488 for the light and heavy sheets respectively (see section III.B for the Supplementary Material[SupMat]). These results show that CFS needs to be treated within the Gutzwiller mean-field theory for correlations to achieve results consistent with experimental observations.
@@ III Conclusion
In the last years there have been a number of studies using various methods to describe the Fermi surface of CeCu2Si2. Experimental measurements have confirmed the existence of a Fermi Surface comprised of two sheets[1990Hunt, 2021Wu], several theoretical and experimental studies have explored the importance of the CFS and electron correlation effects on the overall electronic structure and the superconducting pairing[1993Goremychkin, 2007Ning, 2012Rourke, 2012Willers, 2015Ikeda, 2021Song, 2020Li, 2020Amorese], and new advances have made it possible to perform more accurate and detailed analysis of the intricate coupling and correlation effects in this complex material.
@@
In this study we have presented a systematic analysis of the electronic structure and Fermi surface topology of CeCu2Si2 as a function of the Coulomb interaction strength with a focus on the CFS effect. We have performed calculations using the GWA to show that the resulting Fermi surface topology is highly sensitive to methodology, electron-electron interactions, and CFS treatment. De Haas van Alphen simulations were applied to the Fermi surfaces to analyze the Fermi sheets. Our GWA results calculated using a Hubbard parameter of eV and including CFS in the mean-field theory are found to be in agreement with both experimental measurements as well as the renomalized band theory. More broadly, our results suggest that the CFS plays an important role in driving heavy fermion behavior in the presence of strong correlations in the electronic structure of CeCu2Si2. The qualitative and quantitative aspect of the Fermi surface are also extremely sensitive to small changes in the electron correlation, particularly in the presence of CFS. Based on the analysis of the electronic DOS, this is due to the complexity of the 4f electron behavior present in the heavy Fermi sheet.
@@
Our results highlight the necessity for a thorough treatment of the electronic coupling in order to better understand the nature of the heavy-fermion superconductivity and formation of Cooper pairs. Future studies of the temperature dependence (which has been touched on in Luo et al..[2020Luo]) and possible electron-phonon coupling contribution would aid in forming an even more comprehensive picture of CeCu2Si2.
IV Methods
IV.1 Computational Methods
@@
To account for the strong electronic coupling in CeCu2Si2 we combine standard density functional theory with the GWA method implemented in the CyGutz code [2015Lanata, 2017Lanata]. This method uses the full-potential, linearly augmented plane wave (FPLAPW) DFT of WIEN2k [wien2k] as its basis and implements a combination of the GWA and the rotationally invariant slave boson (RISB) method to account for strong electronic correlations. Relativistic spin-orbit coupling (SOC) effects were included with a k-point grid of and muffin tin radii of 2.50 for the Ce and Cu atoms and 1.93 for the Si atoms, where is the Bohr radius. The cut-off parameter was set to . A generalized gradient approximation (GGA-PBE)[1996Perdew] exchange correlation functional was used throughout all calculations. A double counting correction is done in the fully-localized-limit with the local orbital occupation updated in the outer electron density self-consistent loop, and the Coulomb interaction strength was tested at values from 0 eV to 6.0 eV[SupMat]. We tested the GWA band renormalization implementation with and without CFS effects included beyond the initial DFT. For all calculations where , Hund’s exchange coupling parameter, , was set at 0.7 eV. For comparison, a GGA+DMFT calculation was performed for eV and eV. Results and details of the methodology are given in the Supplementary Materials, sections SII and SIII[SupMat].
IV.2 de Haas van Alphen simulations
@@
Analysis of the Fermi surfaces was done using numerical calculations of the dHvA effect as implemented by Rourke and Julian[2012Rourke]. By applying a magnetic field to the system, oscillations in the magnetic susceptibility can be determined from the changes in the number of occupied Landau levels as a function of the reciprocal magnetic field, [2012Rourke, 2015Jiao]. Then the dHvA frequency can be expressed as
@@ Acknowledgements
The authors thank Chao Cao, Bogyu Jang, Yongxin Yao, E. M. Nica, F. Steglich and G. Zwicknagl for useful discussions. This work was carried out under the auspices of the U.S. Department of Energy (DOE) National Nuclear Security Administration under Contract No. 89233218CNA000001. The Fermi-surface topology analysis work was partly supported by the LANL LDRD Program and the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Science Center. The GWA simulation work was supported by the NNSA Advanced Simulation and Computing Program. It was, in part, supported by the Center for Integrated Nanotechnologies, a DOE BES user facility, in partnership with the LANL Institutional Computing Program for computational resource as well as NERSC.The work at Rice has primarily been supported by the National Science Foundation under Grant No. DMR-2220603, by the Robert A. Welch Foundation Grant No. C-1411 and the Vannevar Bush Faculty Fellowship ONR-VB N00014-23-1-2870. Q. S. acknowledges the hospitality and support by a Ulam Scholarship from the Center for Nonlinear Studies at Los Alamos National Laboratory during the initial stage of the work.